Solve. –6(x + 3) – 10 = 3(–2x – 5) – 13
I keep getting stuck on this one please help?

Answer by funmath(2933)

Solve. –6(x + 3) – 10 = 3(–2x – 5) – 13
-6x-18-10=-6x-15-13
-6x-28=-6x-28
Notice how the left and right side match exactly? That means that this is an identity. It has infinitely many solutions (some teachers say that the solution is all real numbers.) That means that no matter what real number you substitute in for x, the equation will always be equal on both sides.
Happy Calculating!!!